Italian Queue 2014 Prev Next

Jun 19-20

Observer: E. Sani, A. Marchetti, M. Fumana
Telescope Operator: D. Gonzalez-Huerta
Support Astronomer: D. Thompson on call
Instrument: LUCI/LBC/MODS

Summary

At the enclosure opening we have a mostly clear sky (few clouds remaining on the West horizon), no wind and 40% humidity
When scientific operations start we have clear sky with seeing 0.8.
Some difficulties to collimate with LBC for the first target, too low and with still too high illumination from the twilight. We got and error from the blue filter wheel, it happened twice. It was fixed by running again the ob.
Major issues from MODS: stuck of the readout. Reset of the software. Restart of the software. Server not responding --> Restart of the computer. All our, SA and TO efforts failed. Art the end of the night the problem is still unknown. An issue track was send and both the IS (O. Khun) and the PI (R. Pogge) have been notified.
Since we can not run anymore LBC observation due to the Moon, we switch back to LUCI.
The readout noise in LUCI was affecting only the upper right quadrant while acquiring the first target. The effect increased during the acquisition on the last one.
While observing with LUCI the seeing measured on Ks-H2 images is 0.45-0.6
We got the probe error with LUCI, few images are without WFS.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Bellazzini_6Z HVC26.01 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6W HVC28.07 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec PROGRAM COMPLETED
Beltran1_A7 G31.41+0.31 H2, Ks 50min, 8min COMPLETED
Beltran1_A1 G35.20-0.74N H2, Ks 10min, 1.6min partially repeated to update the data from the 1st night

Details

01:53 the TO opens the DOME
  • We run sky flats for program Beltran in H2

02:12 executeLUCIscript.sh dither.flat luci.20140620.0007-0016 GOOD on-flats
  • executeLUCIscript.sh dither.flat luci.20140620.0017-0026 GOOD off-flats ~2000 counts
  • executeLUCIscript.sh dither.flat luci.20140620.0027-0036 GOOD off-flats ~1000 counts
  • executeLUCIscript.sh dither.flat luci.20140620.0037-0046 GOOD off-flats <500 counts
  • executeLUCIscript.sh dither.flat luci.20140620.0057-0056 GOOD off-flats

03:00 We ask the TO to Switch to LBC

03:35 We are authorized to use LBC
  • Slew to Clementini_DDT target
  • dofpia, /x2, /first on copointing target. Few pupils seen on the technical chips, stil light from the twilight. We ask the TO to clean the optics.

03:44 dofpia, /x2, /first on Bellazzini copointing.
  • The script converge but blue images are affected by strong coma. We decide to go ahead since the EL is decreasing

03:58 Run Clementini_DDT_copoint.ob
  • lbcrangebal B=035320 R=035314

04:02 Primary mirro collimation limit reached due to low elevation, we have to abndon the target.
  • the TO cleans the optics

4:08 Slew to Bellazzini_06 Z target
  • dofpia, /x2, /first on Bellazzini copointing. FAILED

04:21 dofpia, /x2 on Bellazzini_06 Z target

04:27 Run 2013B_6_Z_copoint.ob
  • lbcrangebal B=042837 R=042832
  • images affected by coma: lbcb. 043509, 044103, 044200, 044747; lbcr. 043458, 044102, 044151, 044731

04:48 Run dofpia, /X2 on the scientific target

04:54 Run 2013B_6_Z.ob
  • dofpia, /X2
  • still blue image have some coma
  • got an error from the filter wheel #1: timeout error on motion. We call the SA that suggest to just restart the OB.

05:05 We run again dofpia on the scientific target

05:11 Run 6_Z.ob
  • good images, seeing 1.0 in the B channel

05:28 We slew to Bellazzini_6 W target
  • dofpia, /x2 on copointing ob. Coverged at the 2nd time
  • lbcrangebal B=053832 R=05812

05:42 Run 6_W.ob
  • error mooving the blue filter but images lbcb.054441 lbcr.055005 are good

05:51 Run 6_W.ob again

06:10 Run SA_109.ob standard star

06:20 Run SA_109.ob repeat standard star while waiting instruction fro the program to be executed

06:24 We ask the TO to switch to MODS

06:46 We are authorized to use MODS and slew to Elias_Rosa_47_DDT
  • execMODS preima_n1.img to perform 9min integration for pre-imaging.
  • at the end of the readout the process got stuck*. We traied to re-run the script but the sequence failed at the PI name, we modified the string and run the script once more: stuck while cheking the CALIBRATION keyword.
  • There are several efforts to try to recover: the TO restart the software and run sievesahp.pro as a check, it failed at the same point.

07:15 The TO switch off than on the instrument, wake it up and run sieveSnap.pro again. Failed. The TO call the SA.

08.01 After several attempts to reset the system, the SA ask the TO to reset the MODS server and check the connections. We restart everithing, but it failed.

08:25 The SA claim this is something never happened before and it is not straitforward to fix it. The instrument specialist, O. Khun, have to work on it. An issue track is already sent.
  • The SA suggest thus to switch to LUCI

8.48 We are authorized ti use LUCI and slew to the telluric standard for Beltran_01 program.
  • executeLUCIscript.sh FS-148_Ks.img luci.20140620.0057-0061
  • the readout noise is limited and only affecting the upper right quadrant

8.57 executeLUCIscript.sh FS-148_H2.img luci.20140620.0062-0066

9.07 We slew to the scientific target for Beltran A7 with H2 filter
  • executeLUCIscript.sh A7_H2_DJT.img luci.20140620.0067-0086
  • seeing 0.6 and 0.55 on the GCS respectively at the beginning and end of integration.

10:28 executeLUCIscript.sh A7_Ks_DJT.img
  • error from the preset. The TO fix it.

10:32 executeLUCIscript.sh A7_Ks_DJT.img luci.20140620.0067-0086-0106
  • for the 2nd and 4th offset we got the Probe Error Moving. luci.20140620.0098 and luci.20140620.0104 are with no WFS --> no collimation, bad images

10:59 The TO reset the GCS

11:00 executeLUCIscript.sh A7_Ks_DJT.img luci.20140620.0107-0108
  • only for the 2 offsets position where the probe failed in the precedent preset.
  • seeing 0.5 measured in incoming images

11:08 slew to Beltran_01 A1 target, to try to repeat it with limited readout noise.
  • executeLUCIscript.sh A7_H2_DJT.img luci.20140620.0109-0113
  • seeing 0.45 measured in incoming images
  • in the last images the readout noise increased.

11:38 We repeat sky flats in the Ks band
  • executeLUCIscript.sh dither_flat.flat luci.20140620.0116-0225 GOOD flat-off
  • executeLUCIscript.sh dither_flat.flat luci.20140620.0226-0235 GOOD flat-on

12:04 DOME CLOSED
Topic revision: r7 - 23 Jun 2014, EleonoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback